# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v3

@@ 4.2 Language Models with Additional Pretraining on Code
Along with the seminal benchmark HumanEval, Chen et al. (2021b) kick-started the age of LLM for code with Codex, which are GPT-3 checkpoints pretrained on 100B additional code tokens and one of the earliest multi-billion models for code. Following their work, other researchers have also specialized their LLMs on code with additional pretraining. Chowdhery et al. (2022) train PaLM on 7.8B additional code tokens to obtain PaLM-Coder, setting new state-of-the-art on HumanEval and MBPP (Table 1) that are only broken later by its successor PaLM 2-S*, the smallest version of PaLM 2 (Anil et al., 2023) further trained on an undisclosed amount of code. Similarly, Lewkowycz et al. (2022) train PaLM on 38.5B tokens of arXiv papers and mathematical content, while Rozière et al. (2023) train LLaMA 2 (Touvron et al., 2023b) on more than 500B code tokens to acquire Code LLaMA, whose performance on HumanEval surpasses all previous LMs except GPT-4 (Table 1). Liu et al. (2023b) further train Code LLaMA with multi-task finetuning (MFT) to introduce CodeFuse-CodeLLaMA, achieving 74.4 pass@1 on HumanEval and surpassing even the performance of GPT-4 published in OpenAI (2023).
@@ 5.1 Training Dataset for Code
While text data for pretraining language models are often crawled from the web and must undergo meticulous and often aggressive preprocessing (Raffel et al., 2020), code data come naturally as whole documents from public GitHub repositories. Even better, they come with readily available quality indicators such as the count of stars or forks (although Allal et al. (2023) suggest that star count correlates poorly with downstream performance). As a result, many large-scale code pretraining datasets have been introduced, including CodeSearchNet (Husain et al., 2019), CodeParrot (Tunstall et al., 2022), and the Stack (Kocetkov et al., 2022), totaling 20GB, 50GB and 3TB of code documents respectively (Table 2).
@@
Apart from bimodal training and instruction finetuning, another recent trend in constructing code dataset is synthesizing data with powerful models such as ChatGPT. While this method is originally proposed for generating instruction data in natural language (Wang et al., 2023c; Honovich et al., 2023), Gunasekar et al. (2023) take one step further and synthesize 1B tokens of Python textbooks and coding exercises to pretrain a 1.3B model, achieving state-of-the-art results on HumanEval that’s comparable to much larger models trained on significantly larger datasets.
| Dataset | Size (GB) | Files (M) | # PL |
| CodeSearchNet | 20 | 6.5 | 6 |
| The Pile | 95 | 19 | - |
| CodeParrot | 1K | 115 | 30 |
| The Stack | 3136 | 317 | 30 |
| ROOTS | 163 | 15 | 13 |
@@
Date Model Arch. Size Vocab Context PE Preci-sion Atten. Type Parallel Atten. Pre-Norm Flash Atten. Init. from Objectives Dataset Training PL Inst. 2019-12 CuBERT BERT 350M 50K 1024 absolute fp32 MHA - MLM + NSP 9.3B 93B 1 Google 2020-02 CodeBERT RoBERTa 125M 50K 512 absolute fp32 MHA RoBERTa MLM + RTD 20GB 105B 6 Microsoft 2020-09 GraphCode-BERT RoBERTa 125M 50K 640 absolute fp32 MHA CodeBERT MLM + Edge Prediction + Node Alignment 20GB 131B 6 Microsoft 2021-08 SynCoBERT RoBERTa 125M 50K 512 absolute fp32 MHA CodeBERT MLM + IP + AST Edge Prediction + CL 20GB 7B 6 Huawei 2021-10 DISCO BERT 100M 20K 512 absolute fp32 MHA - MLM + Node Type MLM + CL 1.8GB 2 Columbia & IBM 2022-05 Code-MVP RoBERTa 125M 50K 512 absolute fp32 MHA GraphCode-BERT MLM + Type Inference + CL 2GB 39B 1 Huawei 2022-10 SCodeR RoBERTa 125M 51K 1024 absolute fp32 MHA UniXcoder CL 20GB 6 Microsoft 2020-05 GPT-C GPT-2 374M 60K 1024 absolute MHA ✓ - CLM 11B 270B 4 Microsoft 2021-02 CodeGPT GPT-2 124M 50K 1024 absolute fp32 MHA ✓ GPT-2 CLM 2GB 1 Microsoft 2022-02 PolyCoder GPT-NeoX 160M-2.7B 50K 2048 RoPE fp32 MHA ✓ - CLM 254GB 39B 12 CMU 2022-03 CodeGen-Multi(Mono) GPT-J 350M-16.1B 50K 2048 RoPE fp16 MHA ✓ ✓ - CLM 1.6TB(1.8TB)/ 506B(577B) 1T(1.2T) 6(1) Salesforce 2022-04 InCoder GPT-3 6.7B 50K 2048 Cosine fp32 MHA ✓ - Causal Masking 204GB 52B 28 Meta 2022-06 PyCodeGPT GPT-Neo 110M 32K 1024 absolute fp32 MHA ✓ - CLM 96GB 100B 1 Microsoft 2022-07 PanGu-Coder PanGu- 317M-2.6B 42K 1024 absolute MHA ✓ - CLM 147GB 230B 1 Huawei 2023-01 SantaCoder GPT-2 1.1B 49K 2048 absolute fp32 MQA ✓ - FIM 268GB 236B 3 BigCode 2023-03 CodeGeeX PanGu- 13B 52K 2048 absolute fp16 MHA ✓ - CLM 158B 850B 23 Tsinghua 2023-05 StarCoder GPT-2 15.5B 49K 8192 absolute fp32 MQA ✓ ✓ - FIM 815GB 1T 86 BigCode 2023-05 Jam GPT-2 350M 50K 256 absolute bf16 MHA ✓ - CLM 36GB/20B 9B 1 U. Notre Dame 2023-06 Phi-1 GPT-J 1.3B 51K 2048 RoPE fp16 MHA ✓ ✓ ✓ - CLM 7B 53B 1 Microsoft 2023-10 CodeFuse GPT-J 350M-13B 101K 4096 RoPE fp16 MHA ✓ ✓ ✓ - CLM 1.6TB / 1T 40+ Ant Group 2023-10 CodeShell GPT-2 7B 70K 8192 RoPE bf16 GQA ✓ - CLM 500B Peking U. 2020-10 PyMT5 GPT-2 374M 50K 1024+1024 absolute fp16 MHA ✓ - SC 27GB 1 Microsoft 2021-02 Mastropaolo et al. T5 60M 32k 512+512 T5 bf16 MHA ✓ - SC 1GB 1 Universita della Svizzera italiana 2021-02 DOBF 250M 50K 512+512 absolute fp16 MHA - MLM + Deobfuscation 45GB 2 Meta 2021-03 PLBART BART 140M 50K 1024+1024 absolute fp32 MHA - DAE 655GB / 71B 210B 2 UCLA & Columbia 2021-09 CodeT5 T5 60M-220M 32K 512+256 T5 fp16 MHA ✓ - SC + IP + Masked IP + Text2Code + Code2Text 25GB 8 Salesforce 2022-01 SPT-Code BART 262M 80K 512+512 absolute fp16 MHA - NSP + SC + Method Name Prediction 20GB 6 Nanjing U. 2022-02 AlphaCode 300M-41B 8K 1536+768 bf16 MQA - MLM + CLM 715GB 967B 13 DeepMind 2022-06 NatGen T5 220M 32K 512+256 T5 fp16 MHA ✓ CodeT5 Naturalization 26GB 14B 8 Columbia & UC Davis 2022-12 ERNIE-Code mT5 580M 250K 1024+1024 T5 bf16 MHA ✓ mT5 SC + Text2Code + Code2Text + Machine Translation 197B 6 Baidu 2023-05 CodeT5+ T5/GPT-3 220M-16B 50K 2048+2048 absolute fp16 MHA ✓ ✓ CodeGen-mono SC + CLM + CL + Text2Code + Code2Text 52B 9 Salesforce 2020-12 CugLM BERT 51M 50K 128 absolute fp32 MHA - MLM + NSP + CLM 8M 1.2B 2 Peking U. 2022-03 UniXcoder RoBERTa 125M 51K 1024 absolute fp32 MHA - MLM + CLM + SC + CL + Code2Text 20GB+ 839B 6 Microsoft
@@ 5.3 Encoder-Decoders
In NLP, pretrained Transformer encoder-decoders such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020) have also left a notable mark in the past few years’ advancement in language modeling. T5, for example, unifies all textual tasks into a sequence to sequence format and sets new records on GLUE (Wang et al., 2018a) and SuperGLUE (Wang et al., 2019). Compared with encoder-only models, encoder-decoders are naturally more powerful as they can be used for conditional text generation, while their encoder part can always be taken out to perform tasks that require an encoder-only architecture, such as regression (Tay et al., 2023).
@@
Inspired by these advantages of encoder-decoder architecture, many such models have been proposed for code processing. PyMT5 (Clement et al., 2020) and Mastropaolo et al. (2021) replicate the pretraining and multi-task finetuning process of T5 on code corpus, while Ahmad et al. (2021) introduce PLBART, a BART pretrained on 655GB combined data of Java, Pyhton, and natural language. Lachaux et al. (2021) argue that MLM could be too easy a task for programming languages as identifier names often occur multiple times in a single context window, and propose a deobfuscation pretraining objective, where the model is trained to convert obfuscated code back to its original form. Related to this method, we note that meaningful variable names have also been found to have a positive impact on the code generation process of large language models (Chen et al., 2022).
@@
Building on these early works, Wang et al. (2021e) propose CodeT5, which is pretrained alternatively with 1) T5’s original span corruption; 2) identifier tagging (where each token in the code input is tagged as either identifier or non-identifier); 3) masked identifier prediction (a special form of span corruption where all identifiers are masked); and 4) text-to-code & code-to-text generation. Its successor, CodeT5+ (Wang et al., 2023d), take inspiration from UL2 (Tay et al., 2023) and introduce causal language modeling (CLM) into pretraining, along with additional contrastive objectives based on text-code matching.
@@
AlphaCode (Li et al., 2022g) is also trained with multiple objectives, where the encoder is trained with MLM and the decoder is trained with CLM, with architecture modifications such as shallow-encoder & deep-decoder, multi-query attention (Shazeer, 2019), and being much larger than CodeT5 (up to 41B parameters). NatGen (Chakraborty et al., 2022a), on the other hand, is pretrained with a "naturalization" objective similar to deobfuscation: semantically equivalent but unnatural code is generated by predefined operations such as loop transformation, dead code injection, and variable renaming, and the model is pretrained to translate these unnatural code back to its original form. We note that some of these models are built on previous works. For example, NatGen is initialized with CodeT5, while the largest version of CodeT5+ is initialized from a decoder-only model, CodeGen (Nijkamp et al., 2023b).
@@ 5.4 Decoders
After the monumental debut of GPT-3 (Brown et al., 2020) and the discovery of in-context learning, decoder-only Transformer models have become dominant in language modeling (Rae et al., 2021; Hoffmann et al., 2022; Chowdhery et al., 2022; Scao et al., 2022; Touvron et al., 2023a; b, inter alia). Many models similarly pretrained with CLM have also emerged in code processing, such as GPT-C (Svyatkovskiy et al., 2020), CodeGPT (Lu et al., 2021), PolyCoder (Xu et al., 2022), CodeGen (Nijkamp et al., 2023b), PyCodeGPT (Zan et al., 2022), Pangu-Coder (Christopoulou et al., 2022), CodeGeeX (Zheng et al., 2023a), Jam (Su et al., 2023), Phi-1 (Gunasekar et al., 2023), CodeFuse (Di et al., 2023), CodeShell666https://github.com/WisdomShell/codeshell, and DeepSeek Coder777https://github.com/deepseek-ai/DeepSeek-Coder. Of these models, several alternative training objectives have been experimented with, such as MLM and Masked CLM888In their paper, MLM is conducted by replacing tokens in the input with <mask> and predicting it from only the left context, while Masked CLM is performed by adding a <mask> in the input and predict the the next token from it. Both tasks do not change the attention mask patterns of the model. in Pangu-Coder, but are found to underperform compared with CLM-only training. Zan et al. (2022) also propose continual training on sketches, where the model learns to first generate a sketch of a program and then the actual code. Notably, Gunasekar et al. (2023) present Phi-1, a 1.3B small model trained on a dataset of only 7B tokens consisting of 6B tokens from StackOverflow and 1B synthetic data generated by ChatGPT but achieving 50.6 pass@ on HumanEval and 55.5 pass@ on MBPP, comparable to much larger (both in model size and training data size) models such as Code LLaMA or PaLM 2.
| Model | Size | HumanEval | MBPP |
| PolyCoder | 2.7B | 5.6 | - |
| CodeGen-Mono | 16.1B | 29.3 | 35.3 |
| InCoder | 6.7B | 15.2 | 19.4 |
| PyCodeGPT | 110M | 8.3 | - |
| Pangu-Coder | 2.6B | 23.8 | 23.0 |
| SantaCoder | 1.1B | 14.0 | 35.0 |
| CodeGeeX | 13B | 22.9 | 24.4 |
| StarCoder | 15.5B | 33.6 | 52.7 |
| CodeT5+ | 16B | 30.9 | - |
| Phi-1 | 1.3B | 50.6 | 55.5 |
| CodeFuse | 13B | 24.8 | - |
| InstructCodeT5+ | 16B | 35.0 | - |
| WizardCoder | 15.5B | 57.3 | 51.8 |
| Pangu-Coder 2 | 15.5B | 61.6 | - |
| OctoCoder | 15.5B | 46.2 | - |
| CodeFuse-SFT | 13B | 37.1 | - |
| GPT-4 | - | 67.0/82 | - |
| PaLM 2-S* | - | 37.6 | 50.0 |
| Code LLaMA | 34B | 53.7 | 56.2 |
| Phi-1.5 | 1.3B | 41.4 | 43.5 |
@@
Although Christopoulou et al. (2022) report denoising objectives to underperform in decoder-only models, there have been other works that successfully combine denoising or multi-task pretraining with decoder architecture. Incoder (Fried et al., 2023), SantaCoder (Allal et al., 2023), and StarCoder (Li et al., 2023g) are all trained with fill-in-the-middle (FIM) objective, also referred to as causal masking by Fried et al. (2023), which is essentially span corruption (Raffel et al., 2020) adopted to decoder-only architecture. One of the visible advantages of these infilling objectives is that they inject the models with the ability to fill in blanks in the middle of input code at inference time, while CLM allows only for autoregressive generation. As Table 4 shows, however, these objectives also lead to higher performance on downstream tasks when compared with CLM-only models such as CodeGen, although the exact benefits of infilling training remains controversial (Nijkamp et al., 2023a).

### v5

@@ 4.2 Language Models with Additional Pretraining on Code
Along with the seminal benchmark HumanEval, Chen et al. (2021b) kick-started the age of LLM for code with Codex, which are GPT-3 checkpoints pretrained on 100B additional code tokens and one of the earliest multi-billion models for code. Following their work, other researchers have also specialized their LLMs on code with additional pretraining. Chowdhery et al. (2023) train PaLM on 7.8B additional code tokens to obtain PaLM-Coder, setting new state-of-the-art on HumanEval and MBPP (Table 1) that are only broken later by its successor PaLM 2-S*, the smallest version of PaLM 2 (Anil et al., 2023b) further trained on an undisclosed amount of code. Similarly, Lewkowycz et al. (2022) train PaLM on 38.5B tokens of arXiv papers and mathematical content, while Rozière et al. (2023) train LLaMA 2 (Touvron et al., 2023b) on more than 500B code tokens to acquire Code LLaMA, whose performance on HumanEval surpasses all previous LMs except GPT-4 (Table 1). Liu et al. (2023b) further train Code LLaMA with multi-task finetuning (MFT) to introduce CodeFuse-CodeLLaMA, achieving 74.4 pass@1 on HumanEval and surpassing even the performance of GPT-4 published in OpenAI (2023).
@@ 5.1 Training Dataset for Code
While text data for pretraining language models are often crawled from the web and must undergo meticulous and often aggressive preprocessing (Raffel et al., 2020), code data come naturally as whole documents from public GitHub repositories. Even better, they come with readily available quality indicators such as the count of stars or forks (although Allal et al. (2023) suggest that star count correlates poorly with downstream performance). As a result, many large-scale code pretraining datasets have been introduced, including CodeSearchNet (Husain et al., 2019), CodeParrot (Tunstall et al., 2022), and the Stack (Kocetkov et al., 2023), totaling 20GB, 50GB and 3TB of code documents respectively (Table 2).
| Dataset | Size (GB) | Files (M) | # PL |
| CodeSearchNeta | 20 | 6.5 | 6 |
| The Pilebc | 95 | 19 | - |
| CodeParrotd | 1K | 115 | 30 |
| The Stacke | 3136 | 317 | 30 |
| ROOTSf | 163 | 15 | 13 |
| The Stack v2g | 32K | 3K | 619 |
@@
Apart from bimodal training and instruction finetuning, another recent trend in constructing code dataset is synthesizing data with powerful models such as ChatGPT. While this method is originally proposed for generating instruction data in natural language (Wang et al., 2023e; Honovich et al., 2023), Gunasekar et al. (2023) take one step further and synthesize 1B tokens of Python textbooks and coding exercises to pretrain a 1.3B model, achieving state-of-the-art results on HumanEval that’s comparable to much larger models trained on significantly larger datasets.
@@
Date Model Arch. Size Vocab Context PE Preci-sion Atten. Type Parallel Atten. Pre-Norm Flash Atten. Init. from Objectives Dataset Training PL Inst. 2019-12 CuBERT BERT 350M 50K 1024 absolute fp32 MHA - MLM + NSP 9.3B 93B 1 Google 2020-02 CodeBERT RoBERTa 125M 50K 512 absolute fp32 MHA RoBERTa MLM + RTD 20GB 105B 6 Microsoft 2020-09 GraphCode-BERT RoBERTa 125M 50K 640 absolute fp32 MHA CodeBERT MLM + Edge Prediction + Node Alignment 20GB 131B 6 Microsoft 2021-08 SynCoBERT RoBERTa 125M 50K 512 absolute fp32 MHA CodeBERT MLM + IP + AST Edge Prediction + CL 20GB 7B 6 Huawei 2021-10 DISCO BERT 100M 20K 512 absolute fp32 MHA - MLM + Node Type MLM + CL 1.8GB 2 Columbia & IBM 2022-05 Code-MVP RoBERTa 125M 50K 512 absolute fp32 MHA GraphCodeBERT MLM + Type Inference + CL 2GB 39B 1 Huawei 2022-10 SCodeR RoBERTa 125M 51K 1024 absolute fp32 MHA UniXcoder CL 20GB 6 Microsoft 2020-05 GPT-C GPT-2 374M 60K 1024 absolute MHA ✓ - CLM 11B 270B 4 Microsoft 2021-02 CodeGPT GPT-2 124M 50K 1024 absolute fp32 MHA ✓ GPT-2 CLM 2GB 1 Microsoft 2022-02 PolyCoder GPT-NeoX 160M-2.7B 50K 2048 RoPE fp32 MHA ✓ - CLM 254GB 39B 12 CMU 2022-03 CodeGen-Multi(Mono) GPT-J 350M-16.1B 50K 2048 RoPE fp16 MHA ✓ ✓ - CLM 1.6TB(1.8TB)/ 506B(577B) 1T(1.2T) 6(1) Salesforce 2022-04 InCoder GPT-3 6.7B 50K 2048 Cosine fp32 MHA ✓ - Causal Masking 204GB 52B 28 Meta 2022-06 PyCodeGPT GPT-Neo 110M 32K 1024 absolute fp32 MHA ✓ - CLM 96GB 100B 1 Microsoft 2022-07 PanGu-Coder PanGu- 317M-2.6B 42K 1024 absolute MHA ✓ - CLM 147GB 230B 1 Huawei 2023-01 SantaCoder GPT-2 1.1B 49K 2048 absolute fp32 MQA ✓ - FIM 268GB 236B 3 BigCode 2023-03 CodeGeeX PanGu- 13B 52K 2048 absolute fp16 MHA ✓ - CLM 158B 850B 23 Tsinghua 2023-05 StarCoder GPT-2 15.5B 49K 8192 absolute fp32 MQA ✓ ✓ - FIM 815GB 1T 86 BigCode 2023-05 Jam GPT-2 350M 50K 256 absolute bf16 MHA ✓ - CLM 36GB/20B 9B 1 U. Notre Dame 2023-06 Phi-1 GPT-J 1.3B 51K 2048 RoPE fp16 MHA ✓ ✓ ✓ - CLM 7B 53B 1 Microsoft 2023-10 CodeFuse GPT-J 350M-13B 101K 4096 RoPE fp16 MHA ✓ ✓ ✓ - CLM 1.6TB / 1T 40+ Ant Group 2023-10 CodeShell GPT-2 7B 70K 8192 RoPE bf16 GQA ✓ - FIM 100B 500B Peking U. 2020-10 PyMT5 GPT-2 374M 50K 1024+1024 absolute fp16 MHA ✓ - SC 27GB 1 Microsoft 2021-02 Mastropaolo et al. T5 60M 32k 512+512 T5 bf16 MHA ✓ - SC 1GB 1 USI 2021-02 DOBF 250M 50K 512+512 absolute fp16 MHA - MLM + Deobfuscation 45GB 2 Meta 2021-03 PLBART BART 140M 50K 1024+1024 absolute fp32 MHA - DAE 655GB / 71B 210B 2 UCLA & Columbia 2021-09 CodeT5 T5 60M-220M 32K 512+256 T5 fp16 MHA ✓ - SC + IP + Masked IP + TextCode 25GB 8 Salesforce 2022-01 SPT-Code BART 262M 80K 512+512 absolute fp16 MHA - NSP + SC + Method Name Prediction 20GB 6 Nanjing U. 2022-02 AlphaCode 300M-41B 8K 1536+768 bf16 MQA - MLM + CLM 715GB 967B 13 DeepMind 2022-06 NatGen T5 220M 32K 512+256 T5 fp16 MHA ✓ CodeT5 Naturalization 26GB 14B 8 Columbia & UCD 2022-12 ERNIE-Code mT5 580M 250K 1024+1024 T5 bf16 MHA ✓ mT5 SC + TextCode + MT 197B 6 Baidu 2023-05 CodeT5+ T5/GPT-3 220M-16B 50K 2048+2048 absolute fp16 MHA ✓ ✓ CodeGen-mono SC + CLM + CL + TextCode 52B 9 Salesforce 2020-12 CugLM BERT 51M 50K 128 absolute fp32 MHA - MLM + NSP + CLM 8M 1.2B 2 Peking U. 2022-03 UniXcoder RoBERTa 125M 51K 1024 absolute fp32 MHA - MLM + CLM + SC + CL + Code2Text 20GB+ 839B 6 Microsoft
@@ 5.3 Encoder-Decoders
In NLP, pretrained Transformer encoder-decoders such as T5 (Raffel et al., 2020) and BART (Lewis et al., 2020) have also left a notable mark in the past few years’ advancement in language modeling. T5, for example, unifies all textual tasks into a sequence to sequence format and sets new records on GLUE (Wang et al., 2018a) and SuperGLUE (Wang et al., 2019). Compared with encoder-only models, encoder-decoders are naturally more powerful as they can be used for conditional text generation, while their encoder part can always be taken out to perform tasks that require an encoder-only architecture, such as regression (Tay et al., 2023b).
@@
Inspired by these advantages of encoder-decoder architecture, many such models have been proposed for code processing. PyMT5 (Clement et al., 2020) and Mastropaolo et al. (2021) replicate the pretraining and multi-task finetuning process of T5 on code corpus, while Ahmad et al. (2021) introduce PLBART, a BART pretrained on 655GB combined data of Java, Pyhton, and natural language. Lachaux et al. (2021) argue that MLM could be too easy a task for programming languages as identifier names often occur multiple times in a single context window, and propose a deobfuscation pretraining objective, where the model is trained to convert obfuscated code back to its original form. Related to this method, we note that meaningful variable names have also been found to have a positive impact on the code generation process of large language models (Chen et al., 2023c).
@@
Building on these early works, Wang et al. (2021e) propose CodeT5, which is pretrained alternatively with 1) T5’s original span corruption; 2) identifier tagging (where each token in the code input is tagged as either identifier or non-identifier); 3) masked identifier prediction (a special form of span corruption where all identifiers are masked); and 4) text-to-code & code-to-text generation. Its successor, CodeT5+ (Wang et al., 2023f), take inspiration from UL2 (Tay et al., 2023b) and introduce causal language modeling (CLM) into pretraining, along with additional contrastive objectives based on text-code matching.
@@
AlphaCode (Li et al., 2022g) is also trained with multiple objectives, where the encoder is trained with MLM and the decoder is trained with CLM, with architecture modifications such as shallow-encoder & deep-decoder, multi-query attention (Shazeer, 2019), and being much larger than CodeT5 (up to 41B parameters). NatGen (Chakraborty et al., 2022a), on the other hand, is pretrained with a "naturalization" objective similar to deobfuscation: semantically equivalent but unnatural code is generated by predefined operations such as loop transformation, dead code injection, and variable renaming, and the model is pretrained to translate
Model Size HumanEval MBPP PolyCoder 2.7B 5.6 - CodeGen-Mono 16.1B 29.3 35.3 InCoder 6.7B 15.2 19.4 PyCodeGPT 110M 8.3 - Pangu-Coder 2.6B 23.8 23.0 SantaCoder 1.1B 14.0 35.0 CodeGeeX 13B 22.9 24.4 StarCoder 15.5B 33.6 52.7 CodeT5+ 16B 30.9 - Phi-1 1.3B 50.6 55.5 CodeFuse 13B 24.8 - DeepSeek Coder 33B 56.1 66.0 InstructCodeT5+ 16B 35.0 - WizardCoder 15.5B 57.3 51.8 Pangu-Coder 2 15.5B 61.6 - OctoCoder 15.5B 46.2 - CodeFuse 34B 74.4 - DeepSeek Coder-Instruct 33B 79.3 70.0 GPT-4 - 67.0/82 - PaLM 2* S 37.6 50.0 Code LLaMA 34B 53.7 56.2 Phi-1.5 1.3B 41.4 43.5
these unnatural code back to its original form. We note that some of these models are built on previous works. For example, NatGen is initialized with CodeT5, while the largest version of CodeT5+ is initialized from a decoder-only model, CodeGen (Nijkamp et al., 2023b).
@@ 5.4 Decoders
After the monumental debut of GPT-3 (Brown et al., 2020) and the discovery of in-context learning, decoder-only Transformer models have become dominant in language modeling (Rae et al., 2021; Hoffmann et al., 2022; Chowdhery et al., 2023; Scao et al., 2022; Touvron et al., 2023a; b, inter alia). Many models similarly pretrained with CLM have also emerged in code processing, such as GPT-C (Svyatkovskiy et al., 2020), CodeGPT (Lu et al., 2021), PolyCoder (Xu et al., 2022), CodeGen (Nijkamp et al., 2023b), PyCodeGPT (Zan et al., 2022), Pangu-Coder (Christopoulou et al., 2022), CodeGeeX (Zheng et al., 2023a), Jam (Su et al., 2023), Phi-1 (Gunasekar et al., 2023), and CodeFuse (Di et al., 2023). Of these models, several alternative training objectives have been experimented with, such as MLM and Masked CLM666In their paper, MLM is conducted by replacing tokens in the input with <mask> and predicting it from only the left context, while Masked CLM is performed by adding a <mask> in the input and predicting the next token from it. Both tasks do not change the attention mask patterns of the model. in Pangu-Coder, but are found to underperform compared with CLM-only training. Zan et al. (2022) also propose continual training on sketches, where the model learns to first generate a sketch of a program and then the actual code. Notably, Gunasekar et al. (2023) present Phi-1, a 1.3B small model trained on a dataset of only 7B tokens consisting of 6B tokens from StackOverflow and 1B synthetic data generated by ChatGPT but achieving 50.6 pass@ on HumanEval and 55.5 pass@ on MBPP, comparable to much larger (both in model size and training data size) models such as Code LLaMA or PaLM 2.
@@
Although Christopoulou et al. (2022) report denoising objectives to underperform in decoder-only models, there have been other works that successfully combine denoising or multi-task pretraining with decoder architecture. Incoder (Fried et al., 2023), SantaCoder (Allal et al., 2023), StarCoder (Li et al., 2023h), DeepSeek Coder (Guo et al., 2024), and CodeShell (Xie et al., 2024) are trained with fill-in-the-middle (FIM) objective, also referred to as causal masking by Fried et al. (2023), which is essentially span corruption (Raffel et al., 2020) adopted to decoder-only architecture. One of the visible advantages of these infilling objectives is that they inject the models with the ability to fill in blanks in the middle of input code at inference time, while CLM allows only for autoregressive generation. As Table 4 shows, however, these objectives also lead to higher performance on downstream tasks when compared with CLM-only models such as CodeGen, although the exact benefits of infilling training remain controversial (Nijkamp et al., 2023a).